OpenStudy (anonymous):

What is the simplified form of x2-49 (over) x2 - 14x + 49

OpenStudy (anonymous):

x + 7

OpenStudy (anonymous):

\[\frac{x^2-49}{x^2 - 14x + 49}=\frac{(x-7)(x+7)}{(x-7)^2}\]\[=\frac{x+7}{x-7}\]

OpenStudy (anonymous):

that works to

OpenStudy (anonymous):

thank you! :)